# Molecular optomechanics with atomic antennas

### v1

@@ Abstract
A typical surface-enhanced Raman scattering (SERS) system relies on deeply subwavelength field localization in nanoscale plasmonic cavities to enhance both the excitation and emission of Raman-active molecules. Here, we demonstrate that a germanium-vacancy (GeV) defect in diamond can efficiently mediate the excitation process, by acting as a bright atomic antenna. At low temperatures, the GeV’s low dissipation allows it to be efficiently populated by the incident field, resulting in a thousand-fold increase in the efficiency of Raman scattering. Additionally, we show that atomic antenna-enhanced Raman scattering can be distinguished from conventional SERS by tracing the dependence of Stokes intensity on input power, and the pronounced antibunching of the Raman emission.
@@
These developments in alternative platforms have been accompanied by novel theoretical formalisms. In particular, “molecular optomechanics” [11, 12, 13] aims to describe cavity-enhanced Raman scattering by embracing a quantum-mechanical picture of both the optical mode (or modes) and the molecular vibrations. To date, molecular optomechanics has provided insights into various fundamental phenomena such as the optical spring effect and linewidth narrowing [1], dynamical instabilities [14], and correlations between the Stokes and anti-Stokes emission [12, 15]. It has also provided a convenient formalism for describing upconversion of IR photons via Raman transitions [16, 17, 18], and offered insights into the effects of molecular anharmonicities [19, 20], and coherences between vibrations of neighboring molecules [21].
The framework of molecular optomechanics also transparently describes the dual effect of plasmonic cavities in Tip-Enhanced (TERS) and Surface-Enhanced (SERS) Raman Specstroscpy systems [22, 23]. Specifically, cavities serve to (i) localize the field around the molecule, enhancing its driving to a virtual state, and (ii) boost the Stokes and anti-Stokes emission rates by modifying the density of electromagnetic states. In conventional plasmonic SERS, both effects are associated with either one, or several modes of the plasmonic cavity [24, 25], which have linewidths $\kappa$ similar or slightly smaller than the mechanical frequency of the molecular vibrations ($\kappa\lesssim\omega_{b}$). The efficiency of this process can be described by a simple figure of merit $Q/V_{\text{eff}}^{2}$, defined through the effective mode volume $V_{\text{eff}}$ and quality factor $Q$ of the optical mode.
Here we propose an alternative platform to these designs, in the form of a system incorporating an atomic defect — specifically, a germanium vacancy (GeV) in diamond — and a plasmonic nanoparticle (see schematic in Fig. 1(a)). At low temperatures, GeV defects offer two key characteristics ideal for a Raman-enabling antenna: a very low decoherence rate of order $\gamma/2\pi\approx 100$ MHz (about 4 times larger than its lifetime transformed decay rate $\gamma_{0}$), and an unprecedented degree of localisation of the electromagnetic fields. The former means that by acting as an atomic antenna [26], GeV can be far more efficiently excited than any mode of a plasmonic cavity, while the latter ensures that it can drive Raman processes in nearby molecules. At first sight, one might hope to replace both resonant functions of the plasmonic nanoparticle described above with the atomic antenna. However, the atomic linewidth is too narrow to accommodate both the pump and Raman-shifted lines, and it must be partnered with another component, e.g. a gold nanoparticle.
@@
The paper is constructed as follow: in Section I we modify the formalism of molecular optomechanics to explore the differences between the atomic antenna and conventional SERS systems, and compare the two for a set of characteristic parameters in Section II. Finally, in Section III we discuss the effects of the incoherently nonlinear response of the atomic antenna and the resulting characteristics of Raman response.
I Atomic antenna-enhanced Raman scattering
I.1 Coherent dynamics
We consider a Raman dipole $\hat{\mathbf{p}}_{R}$, induced in the molecule positioned at $\mathbf{r}_{m}$ by the electric field $\hat{\mathbf{E}}_{\sigma}$ generated by the excited atomic antenna: $\hat{\mathbf{p}}_{R}=\hat{x}R:\hat{\mathbf{E}}_{\sigma}$. Here $R$ is the Raman tensor, and $\hat{x}$ is the generalized coordinate operator which describes the coordinates of the atoms in the molecule. In the lowest-order approximation [27], we quantize it as a one-dimensional harmonic oscillator, with zero-point fluctuation $Q_{0}=\sqrt{\hbar/(2\omega_{b})}$, frequency $\omega_{b}$, and bosonic operators $\hat{b}/\hat{b}^{\dagger}$ describing the annihilation/creation of the optical phonon. The Raman dipole interacts with the electric field of a particular mode of the plasmonic cavity $\hat{\mathbf{E}}_{a}$, as described by the interaction Hamiltonian [11, 12]
@@ I.1 Coherent dynamics
$=-\frac{1}{2}R_{ij}Q_{0}(\hat{b}+\hat{b}^{\dagger})\Big{(}\hat{\mathbf{E}}_{\sigma}(\mathbf{r}_{m})\Big{)}_{i}\Big{(}\hat{\mathbf{E}}_{a}(\mathbf{r}_{m})\Big{)}_{j}.$ (1)
@@
I.2 Dissipative dynamics
@@
I.3 Linearisation and effective optomechanical coupling
The complete Hamiltonian $\hat{H}_{0}+\hat{H}_{I}+\hat{H}_{\text{coh}}$ can be linearized by realising that in the limit of weak coherent pumping (the so-called Heitler regime) and weak optomechanical interaction, the atomic antenna ($\hat{\sigma}$) will be driven by the incident laser into a predominantly coherent state with amplitude  [^1] In the original papers on molecular optomechanics, we assumed only one plasmonic mode, and attributed the Raman-generated sidebands in the optical mode with the fluctuations in $a$ around its mean.
@@ I.3 Linearisation and effective optomechanical coupling
We revisit this approximation in the last section of the paper, where we explore the Mollow limit of strong pumping of the atomic antenna.
@@
$\hat{H}_{I,\text{lin}}$ $=-\underbrace{\frac{1}{2}RQ_{0}\omega_{\sigma}^{2}\mu_{0}\mathbf{e}_{a}\cdot(\mathbf{G}\cdot\mathbf{d}_{\sigma})^{*}}_{\hbar g_{0,\sigma}}\alpha_{\sigma}\left(\hat{a}+\hat{a}^{\dagger}\right)(\hat{b}+\hat{b}^{\dagger})$
@@
The coefficient $g_{\sigma}=g_{0,\sigma}\alpha_{\sigma}$ is the effective optomechanical coupling, and accounts for the population of the atomic antenna. In conventional cavity optomechanics, these populations can be very large, and compensate for the normally weak $g_{0}$ [28]. Conversely, plasmonic cavities are characterized by very large losses, and cannot be easily populated far above one plasmon ($|\alpha_{a}|=1$). For an atomic antenna, the coherence similarly cannot exceed $|\alpha_{\sigma}|=1$, since the population is by definition limited to unity.
I.4 Emission
@@
$S_{a}(\omega)=\eta_{\text{out}}\omega^{4}\int_{-\infty}^{\infty}\textrm{d}\tau~{}e^{i\omega\tau}\braket{\hat{a}^{\dagger}(\tau)\hat{a}(0)}.$ (10)
@@ I.5 Vibrational populations
The effective dissipation rate of the phonon mode is also modified by these processes as $\Gamma\rightarrow\Gamma_{\text{eff}}=\Gamma+\Gamma_{-}-\Gamma_{+}$. Both effects (the changes in the steady-state phonon population and the linewidth) should be included into (Eq. (I.4)) for the spectrum by substituting $n_{b}\rightarrow n_{b,ss}$ and $\Gamma\rightarrow\Gamma_{\text{eff}}$.
Typical Raman scattering experiments are carried out at room temperature, where the thermal phonon populations $n_{b}^{\text{th}}$ are of the order of $10^{-1}$, and at laser intensities where the modification of the dissipation rate is negligible. In fact, conclusive observation of the change of phonon population due to vibrational pumping (described by the second term on the right-hand side of Eq. (16)), was only reported in [34], by observing a quadratic dependence of the anti-Stokes intensity on pump laser power. Furthermore, observation of a build-up of phonon populations due to the optomechanical heating, represented by the reduction of the effective phonon damping rate $\Gamma_{\text{eff}}$, requires the use of pulsed lasers with large peak powers, to avoid destroying the plasmonic cavities [14].
@@
II Comparison to a conventional plasmonic SERS
@@
In Table 1 we present a comparison of parameters characterising an atomic antenna-enhanced Raman scattering scheme (in a setup schematically represented in Fig. 1), and a conventional plasmonic SERS setup with the same nominal geometry, but without the atomic antenna, where the plasmonic cavity mode mediates both the excitation of the Raman dipole, and its emission. For simplicity, both systems use the same properties of the plasmonic cavity mode (frequency, linewidth and effective mode volume). We note that these parameters are not intended to describe a particular specific plasmonic system, but are characteristic of various nanoparticle-on-mirror and nanoparticle-in-grove platforms [29]. To obtain simple expressions, we also invoke substantial approximations, most significantly (i) assuming that the Green’s function of the system can be approximated by the free-space Green’s function, (ii) taking the coherent driving amplitude $\Omega$ as defined in Eq. (7), without accounting for additional focusing due to the nanoparticle, and (iii) treating the decay rate of the atomic antenna $\gamma$ as independent of the presence of the nanoparticle. We revisit all these approximations in Subsection II.1.
Parameters of the atomic antenna, including the transition dipole moment, are directly taken, or inferred, from Ref. [26] (see Appendix D for the detailed derivation). We assume a shallow location of the defect, 2 nm below the surface of the diamond. We also emphasize that while this work focuses on coherent color centers in diamond, optically coherent defects are observed in several systems, including molecules, rare earth dopants, nanodiamonds, and atomically-thin layered materials such as hexagonal boron nitride [35, 36, 37, 38, 39]. These systems are all suitable candidates for atomic-antenna enhanced Raman spectroscopy.
@@
Parameters $(\omega_{a},\kappa)/2\pi=(498,50)$ THz,
$V_{\text{eff}}=10^{3}~{}\text{nm}^{3}\approx\lambda^{3}/(2\times 10^{5})$,
$\sigma_{\text{ext}}=10^{-13}~{}\text{m}^{2}$ $(\omega_{\sigma},\gamma_{0})/2\pi=(498~{}\text{THz},25~{}\text{MHz})$
$\gamma=4\gamma_{0}$ (at 50 K [26])
$|\mathbf{d}_{\sigma}|\approx 6.7$ D (Appendix D)
$|\mathbf{r}_{\sigma}-\mathbf{r}_{m}|=3$ nm
$RQ^{0}\approx 4\pi\varepsilon_{0}2\times 10^{-30}~{}\text{m}^{3}$$\omega_{b}/2\pi=40$ THz, $\Gamma/2\pi=1$ THz
Electric field $e_{a}\approx 1.4\times 10^{8}$ V/m (Eq. (50)) $e_{\sigma}=\omega_{\sigma}^{2}\mu_{0}G(\mathbf{r}_{\sigma},\mathbf{r}_{m})|\mathbf{d}_{\sigma}|\approx 7.5\times 10^{6}$ V/m
(1 plasmon/exciton)
Optomechanical coupling $\hbar g_{0,a}\approx 1.4\times 10^{-5}$ eV (Eq. (51)) $\hbar g_{0,\sigma}\approx 7\times 10^{-7}$ eV (Eq. (I.3))
$g_{0,a}/2\pi\approx 3.3\times 10^{9}$ Hz $g_{0,\sigma}/2\pi\approx 1.8\times 10^{8}$ Hz
@@
Laser intensity required $I\approx 10^{3}~{}\upmu\text{W}\cdot\upmu\text{m}^{-2}$ $I\approx 6.9\times 10^{-3}~{}\upmu\text{W}\cdot\upmu\text{m}^{-2}$ ($|\alpha_{\sigma}|=1$ in
for 1 plasmon/exciton ($|\alpha_{a}|=1$ in Eq. (48)) in Eq. (8); $60$ deg incidence, $t_{p}=0.46$)
@@
Table 1: Parameters of the systems used to compare the performance of the conventional plasmonic system and atomic antenna. The properties of the plasmonic antenna are shared between the two models, although in the case of the latter system the role of the plasmonic cavity is limited to controlling the Raman emission. Details on the properties of the atomic antenna and the addressed Raman activity are given in Appendix D; equations used to analyse the response of a conventional plasmonic system are given in Appendix E.
@@
The laser intensity which would be required to excite a single coherent plasmon (i.e. to induce a coherent population $|\langle\hat{a}\rangle|=1$) is about $10^{2}~{}\upmu\text{W}\cdot\upmu\text{m}^{-2}$ (see discussion in Appendix E), similar to the maximum CW intensities reported in Ref. [14]. For comparison, the atomic antenna becomes saturated for laser intensities about five orders of magnitude smaller (we discuss the effects of saturation of the atomic dipolar transition in the antenna in Section III). These estimates are consistent with the reported values of power required to saturate other group-IV centers in diamond, e.g. SnV [40].
The single-photon optomechanical coupling in conventional SERS ($g_{0,a}$) is about $20$ times larger than for the atomic antenna ($g_{0,\sigma}$), and both are far smaller than the record values estimated for plasmonic picocavities [2].
@@
We can also estimate the efficiency of down-conversion of the incident to emitted Stokes photons  [^2] A similar question of the conversion efficiency of IR to visible photon was considered by Roelli and co-authors in Ref. [16].. We derive the complete expression for this efficiency in Appendix B, and find that it can be expressed as
@@
where $\eta^{\prime}$ is a prefactor given in Eq. (B). Similarly, we quantify the conversion efficiency for the conventional plasmonic Raman system
@@
Therefore to make meaningful assessments about the resulting contrast in efficiencies, we need to perform a more careful analysis, investigating the effects of the three approximations discussed above.
II.1 Modification of the properties of atomic antenna by a nanoparticle
In the list of parameters given in Table 1, we have neglected the effects of the nanoparticle on the characteristics of the atomic antenna. Here we carry out a more quantitative discussion of these three effects, and their impact on the overall Raman scattering efficiency, supported by numerical solutions implemented in COMSOL Multiphysics software. Details of the numerical model are given in Appendix F. We stress that these calculations are only aimed at establishing the relative corrections to the driving amplitude, decay rate, and field of the atomic antenna, rather than re-assessing all the parameters given in Table 1. Therefore, we keep the remaining values, most critically the frequency and mode volume of the cavity, as constant in the following discussion, meaning that the efficiency of the conventional plasmonic SERS setup remains unchanged.

### v2

@@ Abstract
A typical surface-enhanced Raman scattering (SERS) system relies on deeply subwavelength field localization in nanoscale plasmonic cavities to enhance both the excitation and emission of Raman-active molecules. Here, we demonstrate that a germanium-vacancy (GeV) defect in diamond can efficiently mediate the excitation process, by acting as a bright atomic antenna. At low temperatures, the GeV’s low dissipation allows it to be efficiently populated by the incident field, resulting in a thousand-fold increase in the efficiency of Raman scattering. We show that atomic antenna-enhanced Raman scattering can be distinguished from conventional SERS by tracing the dependence of Stokes intensity on input power.
I Introduction
@@
These developments in alternative platforms have been accompanied by novel theoretical formalisms. In particular, “molecular optomechanics” [11, 12, 13] aims to describe cavity-enhanced Raman scattering by embracing a quantum-mechanical picture of both the optical mode (or modes) and the molecular vibrations. To date, molecular optomechanics has provided insights into various fundamental phenomena such as the optical spring effect and linewidth narrowing and broadening [1], dynamical instabilities [14], and correlations between the Stokes and anti-Stokes emission [12, 15]. It has also provided a convenient formalism for describing upconversion of IR photons via Raman transitions [16, 17, 18], and offered insights into the effects of molecular anharmonicities [19, 20], and coherences between vibrations of neighboring molecules [21].
The framework of molecular optomechanics also transparently describes the dual effect of plasmonic cavities in Tip-Enhanced (TERS) and Surface-Enhanced (SERS) Raman Spectrosocpy systems [22, 23]. Specifically, cavities serve to (i) localize the field around the molecule, enhancing its driving to a virtual state, and (ii) boost the Stokes and anti-Stokes emission rates by modifying the density of electromagnetic states. In conventional plasmonic SERS, both effects are associated with either one, or several modes of the plasmonic cavity [24, 25], which have linewidths $\kappa$ similar to slightly smaller than the mechanical frequency of the molecular vibrations ($\kappa\lesssim\omega_{b}$). The efficiency of this process can be described by a simple figure of merit $Q/V_{\text{eff}}^{2}$, defined through the effective mode volume $V_{\text{eff}}$ and quality factor $Q$ of the optical mode.
Here we propose an alternative platform to these designs in the form of a system incorporating an atomic defect — specifically, a germanium vacancy (GeV) in diamond — and a plasmonic nanoparticle (see schematic in Fig. 1(a)). At low temperatures, GeV defects offer two key characteristics ideal for a Raman-enabling antenna: a very low decoherence rate of order $\gamma/2\pi\approx 100$ MHz (about 4 times larger than its lifetime transformed decay rate $\gamma_{0}$), and an unprecedented degree of localisation of the electromagnetic fields. The former means that by acting as an atomic antenna [26], GeV can be far more efficiently excited than any mode of a plasmonic cavity, while the latter ensures that it can drive Raman processes in nearby molecules. At first sight, one might hope to replace both resonant functions of the plasmonic nanoparticle described above with the atomic antenna. However, the atomic linewidth is too narrow to accommodate both the pump and Raman-shifted lines, and it must be partnered with another component, e.g. a gold nanoparticle.
@@
The paper is constructed as follow: in Section II.1 we modify the formalism of molecular optomechanics to explore the differences between the atomic antenna and conventional SERS systems, and compare the two for a set of characteristic parameters in Section II.2. Finally, in Section II.3 we discuss the effects of the incoherently nonlinear response of the atomic antenna and the resulting characteristics of Raman response.
II Results and discussion
II.1 Atomic antenna-enhanced Raman scattering
A quantum-mechanical description of the off-resonant Raman scattering from a system schematically depicted in Fig. 1(a) can be formulated by explicitly decomposing the optical response of the environment into modes of the plasmonic cavity, and the intrinsically nonlinear atomic antenna. Below we present an abbreviated description of the atomic antenna-enhanced Raman scattering, and direct the Reader to Appendix A for a more complete formulation of this problem.
II.1.1 Coherent dynamics
We consider a Raman dipole $\hat{\mathbf{p}}_{R}$, induced in the molecule positioned at $\mathbf{r}_{m}$ by the electric field $\hat{\mathbf{E}}_{\sigma}$ generated by the excited atomic antenna: $\hat{\mathbf{p}}_{R}=\hat{x}R:\hat{\mathbf{E}}_{\sigma}$. Here $R$ is the Raman tensor, and $\hat{x}$ is the generalized coordinate operator which describes the coordinates of the atoms in the molecule. In the lowest-order approximation [27], we quantize it as a one-dimensional harmonic oscillator, with zero-point fluctuation $Q^{0}=\sqrt{\hbar/(2\omega_{b})}$, frequency $\omega_{b}$, and bosonic operators $\hat{b}/\hat{b}^{\dagger}$ describing the annihilation/creation of the optical phonon. The Raman dipole interacts with the electric field of a particular mode of the plasmonic cavity $\hat{\mathbf{E}}_{a}$, as described by the interaction Hamiltonian [11, 12]
@@ II.1.1 Coherent dynamics
$=-\frac{1}{2}R_{ij}Q^{0}(\hat{b}+\hat{b}^{\dagger})\Big{(}\hat{\mathbf{E}}_{\sigma}(\mathbf{r}_{m})\Big{)}_{i}\Big{(}\hat{\mathbf{E}}_{a}(\mathbf{r}_{m})\Big{)}_{j}.$ (1)
@@
In Appendix A we discuss additional effects that will influence the Raman response of the system, including the coherent driving of the plasmonic mode by incident laser, conventional Raman scattering mediated by that plasmonic mode, and finally its resonant Jaynes-Cummings coupling to the atomic antenna.
II.1.2 Dissipative dynamics
@@
II.1.3 Linearization and effective optomechanical coupling
The complete Hamiltonian $\hat{H}_{0}+\hat{H}_{I}+\hat{H}_{\text{coh}}$ can be linearized by realising that in the limit of weak coherent pumping (the so-called Heitler regime) and weak optomechanical interaction, the atomic antenna ($\hat{\sigma}$) will be driven by the incident laser into a predominantly coherent state with amplitude
@@ II.1.3 Linearization and effective optomechanical coupling
The original papers on molecular optomechanics considered only one plasmonic mode, and attributed the Raman-generated sidebands in the optical mode with the fluctuations in $\hat{a}$ around its mean. We revisit this approximation in the last section of the paper, where we explore the Mollow limit of strong pumping of the atomic antenna.
@@
$\hat{H}_{I,\text{lin}}$ $=-\underbrace{\frac{1}{2}RQ^{0}\omega_{\sigma}^{2}\mu_{0}\mathbf{e}_{a}\cdot(\mathbf{G}\cdot\mathbf{d}_{\sigma})^{*}}_{\hbar g_{0,\sigma}}\alpha_{\sigma}\left(\hat{a}+\hat{a}^{\dagger}\right)(\hat{b}+\hat{b}^{\dagger})$
@@
The coefficient $g_{\sigma}=g_{0,\sigma}\alpha_{\sigma}$ is the effective optomechanical coupling, which accounts for the population of the atomic antenna. In conventional cavity optomechanics, these populations can be very large, and compensate for the normally weak $g_{0}$ [28]. Conversely, plasmonic cavities are characterized by very large losses, and cannot be easily populated far above one plasmon ($|\alpha_{a}|=1$). For an atomic antenna, the coherence similarly cannot exceed $|\alpha_{\sigma}|=2^{-3/2}$ [32].
The exceptionally strong single-photon coupling identified in molecular optomechanics [2, 11] is due to the strong confinement of the electromagnetic field in deeply subwavelength volumes $V_{\text{eff}}$ (via the definition of $g_{0,a}\propto V_{\text{eff}}^{-1}$, see Appendix E). In the hybrid system discussed here, the single-photon optomechanical coupling $g_{0,\sigma}$ is proportional to the field profile $|\mathbf{e}_{a}|\propto V_{\text{eff}}^{-1/2}$.
II.1.4 Emission spectrum
@@
$S_{a}(\omega)=\omega^{4}\int_{-\infty}^{\infty}\textrm{d}\tau~{}e^{i\omega\tau}\braket{\hat{a}^{\dagger}(\tau)\hat{a}(0)}.$ (10)
@@ II.1.5 Vibrational populations
The effective dissipation rate of the phonon mode is also modified by these processes as $\Gamma\rightarrow\Gamma_{\text{eff}}=\Gamma+\Gamma_{-}-\Gamma_{+}$. Both effects (the changes in the steady-state phonon population and the linewidth) should be included into (Eq. (II.1.4)) for the spectrum by substituting $n_{b}\rightarrow n_{b,ss}$ and $\Gamma\rightarrow\Gamma_{\text{eff}}$.
Typical Raman scattering experiments are carried out at room temperature, where the thermal phonon populations $n_{b}^{\text{th}}$ are of the order of $10^{-1}$, and at laser intensities where the modification of the dissipation rate is negligible. In fact, to our best knowledge the first conclusive observation of the change of phonon population due to vibrational pumping (described by the second term on the right-hand side of Eq. (16)), indicated by a quadratic dependence of the anti-Stokes intensity on pump laser power, was reported by Kneipp et al. in Ref. [34]. This study spurred a decade-long effort to understand the different mechanisms that lead to the population of molecular vibrations, and to confirm the central role of vibrational pumping [35, 36, 37, 38]. Observation of a build-up of phonon populations due to the optomechanical backaction, represented by the reduction of the effective phonon damping rate $\Gamma_{\text{eff}}$, required the use of pulsed lasers with large peak powers, to avoid destroying the plasmonic cavities [14].
@@
II.2 Comparison to a conventional plasmonic SERS
@@
In Table 1 we present a comparison of parameters characterising an atomic antenna-enhanced Raman scattering scheme (in a setup schematically represented in Fig. 1), and a conventional plasmonic SERS setup with the same nominal geometry, but without the atomic antenna, where the plasmonic cavity mode mediates both the excitation of the Raman dipole, and its emission (see the schematic of both systems in Fig. 2(a)). For simplicity, both systems use the same properties of the plasmonic cavity mode (frequency, linewidth and effective mode volume). We note that these parameters are not intended to describe a particular specific plasmonic system, but are characteristic of various nanoparticle-on-mirror and nanoparticle-in-grove platforms [29]. To obtain simple expressions, we also invoke substantial approximations, most significantly (i) assuming that the Green’s function of the system can be approximated by the free-space Green’s function, (ii) taking the coherent driving amplitude $\Omega$ as defined in Eq. (7), without accounting for additional focusing due to the nanoparticle, and (iii) treating the decay rate of the atomic antenna $\gamma$ as independent of the presence of the nanoparticle. We revisit all these approximations in Subsection II.2.1.
Parameters of the atomic antenna, including the transition dipole moment, are directly taken, or inferred, from Ref. [26] (see Appendix D for the detailed derivation). We assume a shallow location of the defect, 2 nm below the surface of the diamond. Like other group-IV centres, germanium vacancies GeV exhibit inversion symmetry which decouples the electronic orbitals from the fluctuations of the electric field due to surface charges, protecting the shallow atomic antennas from dephasing. Having said that, we acknowledge that the quoted decoherence rate of $100$ MHz have been measured from bulk defects that are deposited deeper in the crystal. We also emphasize that while this work focuses on coherent color centers in diamond, optically coherent defects are observed in several systems, including molecules, rare earth dopants, nanodiamonds, and atomically-thin layered materials such as hexagonal boron nitride [39, 40, 41, 42, 43]. These systems are all suitable candidates for atomic-antenna enhanced Raman spectroscopy.
@@
Plasmon mode (freq. $\omega_{a}$, linewidth $\kappa$)$/2\pi=(498,50)$ THz,
mode volume $V_{\text{eff}}=10^{3}~{}\text{nm}^{3}\approx\lambda^{3}/(2\times 10^{5})$
ext. cross section $\sigma_{\text{ext}}=10^{-13}~{}\text{m}^{2}$
Atomic antenna (freq. $\omega_{\sigma}$, rad. decay rate $\gamma_{0}$)$/2\pi$
$=(498~{}\text{THz},25~{}\text{MHz})$
tot. decay rate $\gamma=4\gamma_{0}$ (at 50 K [26])
transition dip. mom. $|\mathbf{d}_{\sigma}|\approx 6.7$ D
GeV-mol. distance $|\mathbf{r}_{\sigma}-\mathbf{r}_{m}|=3$ nm
Molecular vibration (freq. $\omega_{b}$, linewidth $\Gamma$)$/2\pi=(40,1)$ THz
Raman strength $RQ^{0}\approx 4\pi\varepsilon_{0}2\times 10^{-30}~{}\text{m}^{3}$
Electric field from $e_{a}\approx 1.4\times 10^{8}$ V/m $e_{\sigma}\approx 7.5\times 10^{6}$ V/m
1 plasmon/exciton (Eq. (43) in SI) (Eq. (38) in SI)
Optomechanical $\hbar g_{0,a}\approx 1.4\times 10^{-5}$ eV $\hbar g_{0,\sigma}\approx 7\times 10^{-7}$ eV (Eq. (II.1.3))
coupling $g_{0,a}/2\pi\approx 3.3\times 10^{9}$ Hz $g_{0,\sigma}/2\pi\approx 1.8\times 10^{8}$ Hz
(Eq. (44) in SI)
(cooperativity) $C_{0,a}\approx 8.6\times 10^{-7}$ $C_{0,\sigma}\approx 1.3\times 10^{-3}$
@@
Laser intensity required $I\approx 10^{3}~{}\upmu\text{W}\cdot\upmu\text{m}^{-2}$ $I\approx 6.9\times 10^{-3}~{}\upmu\text{W}\cdot\upmu\text{m}^{-2}$
for 1 plasmon/exciton
@@
Table 1: Parameters of the systems used to compare the performance of the conventional plasmonic system and atomic antenna. The properties of the plasmonic antenna are shared between the two models, although in the case of the latter system the role of the plasmonic cavity is limited to controlling the Raman emission. Details on the properties of the atomic antenna and the addressed Raman activity are given in Appendix D; equations used to analyse the response of a conventional plasmonic system are given in Appendix E. Laser intensities given in the last row are calculated using Eq. (64), and $|\alpha_{\sigma}|=1$, $60$ deg. incidence and $t_{p}=0.46$ in Eq. (7).
@@
The laser intensity which would be required to excite a single coherent plasmon (i.e. to induce a coherent population $|\langle\hat{a}\rangle|=1$) is about $10^{2}~{}\upmu\text{W}\cdot\upmu\text{m}^{-2}$ (see discussion in Appendix E), similar to the maximum CW intensities reported in Ref. [14]. For comparison, the atomic antenna becomes saturated for laser intensities about five orders of magnitude smaller (we discuss the effects of saturation of the atomic dipolar transition in the antenna in Section II.3). These estimates are consistent with the reported values of power required to saturate other group-IV centers in diamond, e.g. SnV [44].
The single-photon optomechanical coupling in conventional SERS ($g_{0,a}$) is about $20$ times larger than for the atomic antenna ($g_{0,\sigma}$), and both are far smaller than the record values estimated for plasmonic picocavities [2]. Additionally, to compare these couplings with the relevant dissipation and decoherence rates, we estimate the single-photon cooperativities $C_{0,a}=4g_{0}^{2}/(\kappa\Gamma)$ and $C_{0,\sigma}=4g_{0}^{2}/(\gamma\Gamma)$ for the plasmonic and atomic antennas, respectively, and find the latter about 3 orders of magnitude larger. Since both system can only operate below, or near the unitary coherent population of the plasmon or exciton ($|\alpha_{\sigma}|<1$, $|\alpha_{a}|\lesssim 1$), the effective cooperativities $C_{a}=|\alpha_{a}|^{2}C_{0,a}$ and $C_{\sigma}=|\alpha_{\sigma}|^{2}C_{0,\sigma}$ are bounded by similar numbers.
@@
We can also estimate the efficiency of down-conversion of the incident to emitted Stokes photons (a similar question of the conversion efficiency of IR to visible photon was considered by Roelli and co-authors in Ref. [16]). We derive the complete expression for this efficiency in Appendix B, and find that it can be expressed as
@@
where $\eta^{\prime}$ is a prefactor given in Eq. (40). Similarly, we quantify the conversion efficiency for the conventional plasmonic Raman system as
@@
Therefore to make meaningful assessments about the resulting contrast in efficiencies, we need to perform a more careful analysis, investigating the effects of the three approximations listed above.
II.2.1 Modification of the properties of atomic antenna by a nanoparticle
In the list of parameters given in Table 1, we have neglected the effects of the nanoparticle on the characteristics of the atomic antenna. Here we carry out a more quantitative discussion of these three effects, and their impact on the overall Raman scattering efficiency, supported by numerical solutions implemented in COMSOL Multiphysics software. A schematic representation of these systems is shown in Fig. 2(a), and the details of the numerical model are given in Appendix F. We stress that these calculations are only aimed at establishing the relative corrections to the driving amplitude, decay rate, and field of the atomic antenna, rather than re-assessing all the parameters given in Table 1. Therefore, we keep the remaining values, most critically the frequency and mode volume of the cavity, as constant in the following discussion, meaning that the efficiency of the conventional plasmonic SERS setup remains unchanged.